• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name johna

    Log entry time 16:08:08 on March27,2005

    Entry number 142569

    keyword=Shift Summary

    14:00 - Beam went away about 10 minutes ago. MCC calls and says they have to take the beam away for a few minutes to do some tuning to solve the problem.

    14:40 - beam is back. After tuning, MCC was unable to restore the cavity (see ELOG #1263984)